1. Prepare a comparative table of the following concepts:

a. Average Propensity to Consume

b. Average propensity to save
Answer #1

APC:

1.) APC = consumption divided by income = C / Y

APC is slope of consumption function

APC can be greater than one , equal to one or less than one.

APC is always positive , can't be negative

APC falls as income rises.

APS:

1.) APS = saving / income = S/Y

S = Y-C , so APS = 1- APC, or APC + APS= 1

APS is slope of saving function

APS can't be greater than one or equal to one., It's always less than one

APS can be negative

APS rises with rise in income.

As APS rises, APC falls